# CrossMatch: Enhance Semi-Supervised Medical Image Segmentation with Perturbation Strategies and Knowledge Distillation

### v1

@@ Abstract
Semi-supervised learning for medical image segmentation presents a unique challenge of efficiently using limited labeled data while leveraging abundant unlabeled data. Despite advancements, existing methods often do not fully exploit the potential of the unlabeled data for enhancing model robustness and accuracy. In this paper, we introduce CrossMatch, a novel framework that integrates knowledge distillation with dual perturbation strategies—image-level and feature-level—to improve the model’s learning from both labeled and unlabeled data. CrossMatch employs multiple encoders and decoders to generate diverse data streams, which undergo self-knowledge distillation to enhance consistency and reliability of predictions across varied perturbations. Our method significantly surpasses other state-of-the-art techniques in standard benchmarks by effectively minimizing the gap between training on labeled and unlabeled data and improving edge accuracy and generalization in medical image segmentation. The efficacy of CrossMatch is demonstrated through extensive experimental validations, showing remarkable performance improvements without increasing computational costs. Code for this implementation is made available at https://github.com/AiEson/CrossMatch.git.
@@ 1 Introduction
Semantic segmentation, as a precise classification technique at the pixel level, plays a vital role in the field of medical image analysis. Especially when dealing with complex three-dimensional CT and MRI data, although fully supervised learning methods can achieve high-precision segmentation results, their application is severely limited by the high cost of manual annotation and the complexity of operation. In order to overcome this bottleneck, semi-supervised medical image segmentation methods have emerged and demonstrate great potential [1]. The core of these approaches lies in the effective combination of a small amount of annotated data and a large amount of unlabeled data, aiming to reduce the high cost of annotation and achieve accuracy segmentation while promoting the widespread application in clinical and other scenarios.
@@
The main challenge in semi-supervised learning (SSL) is how to effectively exploit the potential of unlabeled data. Recent research has shifted from relying on adversarial training mechanisms based on Generateve Adversarial Networks (GANs) [2, 3] to incorporating various methods including consistency regularization and self-training [4, 5, 6, 7, 8]. In particular, collaborative teaching and mutual learning paradigms [9, 10, 11, 12, 13] have proven to be highly promising strategies, often involving the parallel training of two models. Knowledge distillation strategies have also been widely employed to optimize model structures, enabling efficient training and good performance by simplifying models.
@@
In handling unlabelled image data, the application of both image-level and feature-level perturbations has become a common strategy. Image-level perturbations, such as random rotations, scaling, flipping, and color adjustments, enhance model robustness to input variations through controlled deformations and modifications of the input images. Moreover, more complex image-level perturbations like CutMix [14] and MixUp [15] create new training samples by blending regions between images and combining them at the pixel level, thus simulating a more diverse data distribution and further improving the model’s generalization to unseen data. Feature-level perturbations, particularly those applied to features extracted by the Encoder, have not been fully explored and hold substantial potential. This approach introduces weak to strong feature perturbations during the Decoder decoding process, utilizing the model’s prediction consistency under various perturbation conditions to train the model, which ensures stability in performance when the model faces the same image segmentation tasks. For example, feature-level perturbations can be achieved by adding random noise, applying various types of Dropout, etc. [8, 16]. These perturbations not only simulate potential variations in the data but also promote generalization in the model’s deep feature abstraction and decoding processes, thereby achieving more accurate and robust predictive performance on unlabeled data.
@@
Knowledge Distillation(KD) has demonstrated significant potential in semi-supervised learning for medical image segmentation [17, 13]. Typically, KD involves a pre-trained teacher model and a student model that needs to learn. However, Self-KD methods [18, 19] primarily rely on soft labels generated within a single model to guide the training process, instead of depending on traditional hard labels or an additional teacher model. These methods use the model’s self-generated predictions during training as guidance, refining the model’s feature extraction and classification capabilities through iterative processes. This self-teaching method not only reduces dependence on costly manually annotated data but also significantly enhances the model’s adaptability and prediction accuracy on unlabeled data. Self-KD promotes deeper feature learning and more stable model behavior by reinforcing the model’s reliance on its own predictions. Particularly for medical imaging data, this strategy effectively improves model robustness and accuracy when dealing with highly variable and individually distinct medical images.
@@
Inspired by Self-KD and image perturbation, we have designed an innovative self-training consistency regularization framework called CrossMatch for semi-supervised medical image segmentation. This framework employes a range of image-level and feature-level perturbations from weak to strong and explores the potential of unlabeled data through a more systematic and in-depth approach. Specifically, CrossMatch applies two types of image-level and two types of feature-level perturbations to unlabeled data to create four distinct data streams. These data streams vary in accuracy of output prediction depending on the degree of perturbation to which they are subjected, where the stronger streams guide the weaker ones. In this process, image-level perturbations are implemented as applications of different encoders, while feature-level perturbations are used to generate varied outputs for the same decoder. Through these perturbations, CrossMatch engages in internal knowledge distillation by leveraging the model’s consistency across different perturbation intensities, which not only optimizes the model’s learning from unlabeled data but also enhances its generalization capability. CrossMatch ensures the stability and accuracy of model outputs, thereby exhibiting superior performance in applications requiring high precision, such as medical image segmentation.
@@
    Experimental results on two benchmark datasets demonstrate that CrossMatch achieves significant performance improvements compared to previous state-of-the-art methods.
@@ 2.1 Semi-Supervised Learning
In the field of SSL, a key challenge is designing effective supervision signals for unlabeled data. Currently, there are two main strategies to address this issue: entropy minimization [20, 21, 22, 23] and consistency regularization [6, 24, 25, 8]. Entropy minimization, popular for its simplicity, involves automatically assigning pseudo-labels to unlabeled data and using them for retraining along with labeled data. And consistency regularization is based on the assumption that a model’s predictions for the same unlabeled sample under different perturbations should remain consistent. For example, FixMatch [6] combines the advantages of entropy minimization and consistency regularization to apply strong perturbations to unlabeled images and use the predictions from their weakly perturbed versions to guide model training. Advanced methods like FreeMatch [26] further refine this strategy, providing rigorous mathematical justification for its motivation and using thresholds to filter out low-confidence labels, thereby enhancing the model’s accuracy and reliability. Our CrossMatch draws on the basic framework of FixMatch, without any bells and whistles. It only uses the most common way to verify the theoretical effectiveness of this method, and also demonstrates its important value in practical applications.
@@ 2.2 Semi-Supervised Semantic Segmentation
Semi-supervised learning based methods have achieved exciting results in classification task, of which several works have been further developed for semantic segmentation. A popular class of methods [9, 27, 28] is based on the Mean Teacher [29] setting. For instance, UA-MT [27] introduces a self-aware model of uncertainty to design thresholds to filter out uncertain regions between teachers and students to get more meaningful and reliable predictions. BCP [28] notes that in semi-supervised learning, the distributions learned in labeled and unlabeled data are not consistent, and proposes a symmetric approach to use both kinds of data so as to maintain the consistency between the two distributions, thus allowing the model to learn common features. CAML [9] pays further attention to the potential of labeled data and proposes Correlation Aware Mutual Learning framework to utilize labeled data to guide the extraction of information from unlabeled data. CPS [4] utilizes a cross-teacher module to simultaneously reduce the coupling among peer networks and the error accumulation between teacher and student networks.
@@
Another mainstream class of semi-supervised segmentation methods is based on the idea of co-training. The networks learn together and transfer knowledge to each other [10, 13]. To transfer knowledge efficiently between networks, knowledge distillation is also a common strategy in semi-supervised semantic segmentation [13]. Besides, some method uses pseudo segmentation map obtained from one network to supervise the other one [30]. MC-Net [11] and MC-Net+ [12] use a shared encoder for feature extraction, and then feed the features into multiple decoders with the same structure but different parameters to get multiple outputs. All these methods require multiple networks, encoders or decoders for training.
@@
Methods based on self-training have begun to evolve rapidly since FixMatch [6] introducing consistency regularisation to self-training, and FixMatch has gradually become the baseline for many methods. DTC [31] uses a dual-task deep network to jointly predict pixel segmentation maps and geometrically-aware level-set representations of a target by introducing dual-task consistency regularisation between level-set derived segmentation maps and directly predicted segmentation maps for labelled and unlabelled data. SASSNet [32] introduces a multi-task deep network that jointly predicts semantic segmentation and symbolic distance maps (SDM) of object surfaces, while introducing adversarial loss in order to capture shape-aware features. URPC [33] enhances pyramid-consistent regularisation using multi-scale uncertainty correction for more efficient semi-supervised medical image segmentation. SS-Net [34] addresses the challenges of semi-supervised medical image segmentation by simultaneously exploring pixel-level smoothness and inter-class separation. UniMatch [8] achieves better segmentation results by consistency regularisation using multiple strongly augmented branches and a dual-stream perturbation feature perturbation. Our CrossMatch also follows this single-stage framework, i.e., there is only one model in our approach. Unlike the above works, our CrossMatch introduces Self-KD and feature perturbation into semi-supervised medical image segmentation, achieving efficient self-knowledge transfer under a broader perturbation space.
@@ 3.1 Preliminaries
Figure 1: Comparison of different types of KD and SSL methods. (a) Traditional KD that requires pre-training of the teacher model. (b) Self-KD based on data augmentation. (c) Mean Teacher. (d) FixMatch.
@@ 3.2 Knowledge Distillation
where $p^{w_{i}}$ and $p^{w_{j}}$ represent the probability distribution outputs by the teacher and student models for unlabeled samples, respectively. Here, $\sigma(\cdot)$ denotes the softmax function, which transforms logits into normalized probability distributions, and $T$ is a hyperparameter known as the temperature coefficient. This coefficient smooths the probability distributions, allowing the student model not only to focus on the most likely predicted category but also to learn more about the relative information between categories from the teacher model. Thus, within the framework of knowledge distillation, the KL divergence loss function $\mathcal{L}_{kd}^{KL}$ aims to minimize the difference between the probability distributions of the teacher model $w_{i}$ and the student model $w_{j}$ that after softmax processing and temperature reduction. Through this way, the student model can emulate the teacher model’s ’soft’ output predictions, thereby facilitating effective transfer of complex and high-quality knowledge.
DMD [13] delves into knowledge distillation methods specifically for semi-supervised medical image segmentation, and proposes to use Dice loss as an alternative to KL divergence loss. This approach effectively addresses the common issue of foreground and background class imbalance in segmentation tasks. Compared to KL divergence loss, Dice loss can more aptly handle such imbalances, thereby enhancing the model’s segmentation performance,
@@ 3.3 Feature Pertubation
Although image-level perturbations have widely used in numerous methods, their performance in semi-supervised image segmentation tasks highly depends on how researchers meticulously tailor perturbation schemes for specific datasets to ensure an appropriate perturbation space is constructed. This process often involves a high demand for expert knowledge and trial-and-error costs, especially in the field of medical image processing, where finding suitable perturbation strategies can become one of the main challenges demanding significant effort.
To mitigate the aforementioned issues, the literature [8, 16] suggests perturbing the high-dimensional features of $x^{w}$ at the bottleneck section of the segmentation network by using different levels of perturbation to create varied feed-forward flows. Segmentation models typically employ an encoder-decoder structure, where $e$ denotes the encoder and $d$ denotes the decoder. For FixMatch, the weak perturbation feed-forward flow for an unlabeled sample $x^{u}$ can be represented as:
@@
$\begin{matrix}x^{w}\to e\to\mathcal{P}^{r}\to d\to p^{w}_{r}\\ x^{s}\to e\to\mathcal{P}^{r}\to d\to p^{s}_{r}\end{matrix}.$ (7)
@@
where $\mathrm{H}(\cdot)$ denotes the entropy minimizing the discrepancy between two probability distributions. $\mathbbm{1}(\cdot>\tau)$ is the indicator function for confidence-based thresholding with the threshold $\tau$. $B_{u}$ is the batch size for unlabeled data.
@@ 3.5 CrossMatch
Fig. 1 demonstrates various KD and SSL methods, revealing a high similarity between them, thus prompting the idea of integrating knowledge distillation into SSL tasks. The purpose of knowledge distillation is to transfer more accurate knowledge to another network model. Self-KD represents a unique distillation mode where the student model learns from knowledge generated from its own outputs, typically involving the backpropagation of deep information to guide the training of earlier layers. This approach incorporates image-level perturbations to achieve varying capabilities, as depicted in Fig. 1 (b), a process very similar to that in Fig. 1 (c).
@@
${\color[rgb]{0,0,0}\mathcal{L}_{tkd}}=\frac{1}{B_{u}}\sum\mathbbm{1}\left(\max\left(p^{w}_{n}\right)\geq\tau\right)\sum_{i}\sum_{j}\mathrm{H}\left(p^{w}_{n},p^{i}_{j}\right).$ (10)
Observing that a decoder outputs two segmentation results with varying degrees of perturbation, we can facilitate mutual distillation between these outputs. Specifically, we consider $p^{w}_{w}$ and $p^{w}_{s}$ as teaching assistants, each imparting knowledge to $p^{s}_{w}$ and $p^{s}_{s}$, respectively. These teaching assistants are relative to the same decoder, hence this is referred to as decoder distillation loss:
${\color[rgb]{0,0,0}\mathcal{L}_{dkd}}=\frac{1}{B_{u}}\sum\mathbbm{1}\left(\max\left(p^{w}_{j}\right)\geq\tau\right)\sum_{j}\mathrm{H}\left(p^{w}_{j},p^{s}_{j}\right).$ (11)
@@
where $\mathcal{L}_{sup}$ consists of Dice and CrossEntropy losses, and $\mathcal{L}_{ip}$ denotes the supervision of $p^{w}_{n}$ over $p^{s_{1}}$ and $p^{s_{2}}$ as shown in Fig. 3, which involves calculating Dice for the two strongly perturbed unlabeled predictions and averaging them. $\eta$ is used to balance the proportions between the two distillation loss.

### v2

@@ Abstract
Semi-supervised learning for medical image segmentation presents a unique challenge of efficiently using limited labeled data while leveraging abundant unlabeled data. Despite advancements, existing methods often do not fully exploit the potential of the unlabeled data for enhancing model robustness and accuracy. In this paper, we introduce CrossMatch, a novel framework that integrates knowledge distillation with dual perturbation strategies, image-level and feature-level, to improve the model’s learning from both labeled and unlabeled data. CrossMatch employs multiple encoders and decoders to generate diverse data streams, which undergo self-knowledge distillation to enhance the consistency and reliability of predictions across varied perturbations. Our method significantly surpasses other state-of-the-art techniques in standard benchmarks by effectively minimizing the gap between training on labeled and unlabeled data and improving edge accuracy and generalization in medical image segmentation. The efficacy of CrossMatch is demonstrated through extensive experimental validations, showing remarkable performance improvements without increasing computational costs. Code for this implementation is made available at https://github.com/AiEson/CrossMatch.git.
@@ 1 Introduction
Semantic segmentation, as a precise classification technique at the pixel level, plays a vital role in the field of medical image analysis. Especially when dealing with complex three-dimensional CT and MRI data, although fully supervised learning methods can achieve high-precision segmentation results, their applications are severely limited by the high cost of manual annotation and the complexity of operation. In order to overcome this bottleneck, semi-supervised medical image segmentation methods have emerged, and demonstrated great potential [1]. The core of these approaches lies in the effective combination of a small amount of annotated data and a large amount of unlabeled data, aiming to reduce the high cost of annotation and achieve accurate segmentation while promoting widespread application in clinical and other scenarios.
@@
The main challenge in semi-supervised learning (SSL) is how to exploit the potential of unlabeled data effectively. Recent research has shifted from relying on adversarial training mechanisms based on Generative Adversarial Networks (GANs) [2, 3] to incorporating various methods, including consistency regularization and self-training [4, 5, 6, 7, 8]. In particular, collaborative teaching and mutual learning paradigms [9, 10, 11, 12, 13] have proven to be highly promising strategies, often involving the parallel training of two models. Knowledge distillation strategies have also been widely employed to optimize model structures, enabling efficient training and good performance by simplifying models.
@@
In handling unlabeled image data, the application of both image-level and feature-level perturbations has become a common strategy. Image-level perturbations, such as random rotations, scaling, flipping and color adjustments, enhance model robustness to input variations through controlled deformations and modifications of the input images. Moreover, more complex image-level perturbations like CutMix [14] and MixUp [15] create new training samples by blending regions between images and combining them at the pixel level, thus simulating a more diverse data distribution and further improving the model’s generalization to unseen data. Feature-level perturbations, particularly those applied to features extracted by the Encoder, have not been fully explored and hold substantial potential. This approach introduces weak to strong feature perturbations during the Decoder decoding process, utilizing the model’s prediction consistency under various perturbation conditions to train the model, which ensures stability in performance when the model faces the same image segmentation tasks. For example, feature-level perturbations can be achieved by adding random noise, applying various types of Dropout, etc. [8, 16]. These perturbations not only simulate potential variations in the data but also promote generalization in the model’s deep feature abstraction and decoding process, thereby achieving more accurate and robust predictive performance on unlabeled data.
@@
Knowledge Distillation (KD) has demonstrated significant potential in semi-supervised learning for medical image segmentation [17, 13]. Typically, KD involves a pre-trained teacher model and a student model that needs to be learned. However, Self-KD methods [18, 19] primarily rely on soft labels generated within a single model to guide the training process instead of depending on traditional hard labels or an additional teacher model. These methods use the model’s self-generated predictions during training as guidance, refining the model’s feature extraction and classification capabilities through iterative processes. This self-teaching method not only reduces dependence on costly manually annotated data but also significantly enhances the model’s adaptability and prediction accuracy on unlabeled data. Self-KD promotes deeper feature learning and more stable model behaviour by reinforcing the model’s reliance on its own predictions. Particularly for medical imaging data, this strategy effectively improves model robustness and accuracy when dealing with highly variable and individually distinct medical images.
@@
Inspired by Self-KD and image perturbation, we have designed an innovative self-training consistency regularization framework called CrossMatch for semi-supervised medical image segmentation. This framework employs a range of image-level and feature-level perturbations from weak to strong and explores the potential of unlabeled data through a more systematic and in-depth approach. Specifically, CrossMatch applies two types of image-level and two types of feature-level perturbations to unlabeled data to create four distinct data streams. These data streams vary in accuracy of output prediction depending on the degree of perturbation to which they are subjected, where the stronger streams guide the weaker ones. In this process, image-level perturbations are implemented as applications of different encoders, while feature-level perturbations are used to generate varied outputs for the same decoder. Through these perturbations, CrossMatch engages in internal knowledge distillation by leveraging the model’s consistency across different perturbation intensities, which not only optimizes the model’s learning from unlabeled data but also enhances its generalization capability. CrossMatch ensures the stability and accuracy of model outputs, thereby exhibiting superior performance in applications requiring high precision, such as medical image segmentation.
@@
    Experimental results on four benchmark datasets demonstrate that CrossMatch achieves significant performance improvements compared to previous state-of-the-art methods.
@@ 2.1 Semi-Supervised Learning
In the field of SSL, a key challenge is designing effective supervision signals for unlabeled data. Currently, there are two main strategies to address this issue: entropy minimization [20, 21, 22, 23] and consistency regularization [6, 24, 25, 8]. Entropy minimization, a concept utilized in semi-supervised learning settings, involves measuring the difference in entropy between the outputs of teacher and student models. By minimizing this difference, we ensure that the student model’s predictions are confident and closely aligned with those of the teacher model. This strategy aids in stabilizing the learning process by fostering more reliable and consistent predictions from the student model. Incorporating this into the broader context of entropy minimization and consistency regularization, we can see how these strategies synergize. Entropy minimization is appreciated for its straightforward approach of automatically assigning pseudo-labels to unlabeled data. These pseudo-labels are then used to retrain the model alongside labeled data, enhancing the model’s ability to generalize from limited labeled information. Consistency regularization complements this by ensuring that a model’s predictions for the same unlabeled sample remain consistent across different perturbations. This consistency is crucial, as it helps in reducing overfitting and improving the robustness of the model under varying conditions. For example, FixMatch [6] combines the advantages of entropy minimization and consistency regularization to apply strong perturbations to unlabeled images and use the predictions from their weakly perturbed versions to guide model training. Advanced methods like FreeMatch [26] further refine this strategy, providing rigorous mathematical justification for its motivation and using thresholds to filter out low-confidence labels, thereby enhancing the model’s accuracy and reliability.
Our CrossMatch draws on the basic framework of FixMatch without any bells and whistles. It only uses the most common way to verify the theoretical effectiveness of this method and also demonstrates its important value in practical applications.
@@ 2.2 Semi-Supervised Semantic Segmentation
Semi-supervised learning based methods have achieved exciting results in classification tasks, of which several works have been further developed for semantic segmentation. A popular class of methods [9, 27, 28] is based on the Mean Teacher [29] setting. For instance, UA-MT [27] introduces a self-aware model of uncertainty to design thresholds to filter out uncertain regions between teachers and students to get more meaningful and reliable predictions. BCP [28] notes that in semi-supervised learning, the distributions learned in labeled and unlabeled data are not consistent and proposes a symmetric approach to use both kinds of data so as to maintain the consistency between the two distributions, thus allowing the model to learn common features. CAML [9] pays further attention to the potential of labeled data and proposes a Correlation Aware Mutual Learning framework to utilize labeled data to guide the extraction of information from unlabeled data. CPS [4] utilizes a cross-teacher module to simultaneously reduce the coupling among peer networks and the error accumulation between teacher and student networks.
@@
Another mainstream class of semi-supervised segmentation methods is based on the idea of co-training. The networks learn together and transfer knowledge to each other [10, 13]. To transfer knowledge efficiently between networks, knowledge distillation is also a common strategy in semi-supervised semantic segmentation [13]. Besides, some method uses pseudo segmentation maps obtained from one network to supervise the other one [30]. MC-Net [11] and MC-Net+ [12] use a shared encoder for feature extraction and then feed the features into multiple decoders with the same structure but different parameters to get multiple outputs. All these methods require multiple networks, encoders or decoders for training.
@@
Methods based on self-training have begun to evolve rapidly since FixMatch [6] introduced consistency regularization to self-training, and FixMatch has gradually become the baseline for many methods. DTC [31] uses a dual-task deep network to jointly predict pixel segmentation maps and geometrically-aware level-set representations of a target by introducing dual-task consistency regularization between level-set derived segmentation maps and directly predicted segmentation maps for labeled and unlabeled data. SASSNet [32] introduces a multi-task deep network that jointly predicts semantic segmentation and symbolic distance maps (SDM) of object surfaces while introducing adversarial loss in order to capture shape-aware features. URPC [33] enhances pyramid-consistent regularization using multi-scale uncertainty correction for more efficient semi-supervised medical image segmentation. SS-Net [34] addresses the challenges of semi-supervised medical image segmentation by simultaneously exploring pixel-level smoothness and inter-class separation. UniMatch [8] achieves better segmentation results by consistency regularization using multiple strongly augmented branches and a dual-stream perturbation feature perturbation. Our CrossMatch also follows this single-stage framework, i.e., there is only one model in our approach. Unlike the above works, our CrossMatch introduces Self-KD and feature perturbation into semi-supervised medical image segmentation, achieving efficient self-knowledge transfer under a broader perturbation space.
@@ 3.1 Preliminaries
Figure 1: Comparison of different types of KD and SSL methods. (a) Traditional KD requires pre-training of the teacher model. (b) Self-KD based on data augmentation. (c) Mean Teacher. (d) FixMatch.
@@ 3.2 Knowledge Distillation
where $p^{w_{i}}$ and $p^{w_{j}}$ represent the probability distribution outputs by the teacher and student models for unlabeled samples, respectively. Here, $\sigma(\cdot)$ denotes the softmax function, which transforms logits into normalized probability distributions, and $T$ is a hyperparameter known as the temperature coefficient. This is a scaling factor used in the softmax function during knowledge distillation to control the smoothness of output probabilities. A higher temperature results in softer probability distributions, facilitating the transfer of richer information between the teacher and student models. Thus, within the framework of knowledge distillation, the KL divergence loss function $\mathcal{L}_{kd}^{KL}$ aims to minimize the difference between the probability distributions of the teacher model $w_{i}$ and the student model $w_{j}$ that after softmax processing and temperature reduction. In this way, the student model can emulate the teacher model’s ’soft’ output predictions, thereby facilitating the effective transfer of complex and high-quality knowledge.
DMD [13] delves into knowledge distillation methods specifically for semi-supervised medical image segmentation and proposes to use Dice loss as an alternative to KL divergence loss. This approach effectively addresses the common issue of foreground and background class imbalances in segmentation tasks. Compared to KL divergence loss, Dice loss can more aptly handle such imbalances, thereby enhancing the model’s segmentation performance,
@@ 3.3 Feature Pertubation
Although image-level perturbations have been widely used in numerous methods, their performance in semi-supervised image segmentation tasks highly depends on how researchers meticulously tailor perturbation schemes for specific datasets to ensure an appropriate perturbation space is constructed. This process often involves a high demand for expert knowledge and trial-and-error costs, especially in the field of medical image processing, where finding suitable perturbation strategies can become one of the main challenges demanding significant effort.
To mitigate the issues mentioned above, the literature [8, 16] suggests perturbing the high-dimensional features of $x^{w}$ at the bottleneck section of the segmentation network by using different levels of perturbation to create varied feed-forward flows. Segmentation models typically employ an encoder-decoder structure, where $e$ denotes the encoder and $d$ denotes the decoder. For FixMatch, the weak perturbation feed-forward flow for an unlabeled sample $x^{u}$ can be represented as:
@@
$\left\{\begin{array}[]{l}x^{w}\to e\to\mathcal{P}^{r}\to d\to p^{w}_{r}\\ x^{s}\to e\to\mathcal{P}^{r}\to d\to p^{s}_{r}.\end{array}\right.$ (7)
@@
where $\mathrm{H}(\cdot)$ denotes the entropy minimizing the discrepancy between two probability distributions. $\mathbbm{1}(\cdot>\tau)$ is the indicator function for confidence-based thresholding with the threshold $\tau$. $\max(\cdot)$ represents taking the maximum value along the channel dimension to obtain the confidence map. $B_{u}$ is the batch size for unlabeled data.
@@ 3.5 CrossMatch
Fig. 1 demonstrates various KD and SSL methods, revealing a high similarity between them, thus prompting the idea of integrating knowledge distillation into SSL tasks. The purpose of knowledge distillation is to transfer more accurate knowledge to another network model. Self-KD represents a unique distillation mode where the student model learns from knowledge generated from its outputs, typically involving the backpropagation of deep information to guide the training of earlier layers. This approach incorporates image-level perturbations to achieve varying capabilities, as depicted in Fig. 1 (b), a process very similar to that in Fig. 1 (c).
@@
${\mathcal{L}_{tkd}}=\frac{1}{B_{u}}\sum\mathbbm{1}\left(\max\left(p^{w}_{n}\right)\geq\tau\right)\sum_{i}\sum_{j}\mathrm{H}\left(p^{w}_{n},p^{i}_{j}\right).$ (10)
Observing that a decoder outputs two segmentation results with varying degrees of perturbation, we can facilitate mutual distillation between these outputs. Specifically, we consider $p^{w}_{w}$ and $p^{w}_{s}$ as teaching assistants, each imparting knowledge to $p^{s}_{w}$ and $p^{s}_{s}$, respectively. These teaching assistants are relative to the same decoder hence this is referred to as decoder distillation loss:
${\mathcal{L}_{dkd}}=\frac{1}{B_{u}}\sum\mathbbm{1}\left(\max\left(p^{w}_{j}\right)\geq\tau\right)\sum_{j}\mathrm{H}\left(p^{w}_{j},p^{s}_{j}\right).$ (11)
@@
where $\mathcal{L}_{sup}$ consists of Dice and CrossEntropy losses, and $\mathcal{L}_{ip}$ denotes the supervision of $p^{w}_{n}$ over $p^{s_{1}}$ and $p^{s_{2}}$ as shown in Fig. 3, which involves calculating Dice for the two strongly perturbed unlabeled predictions and averaging them. $\eta$ is used to balance the proportions between the two distillation losses.